OpenStudy (anonymous):

integrate 1/9+sqrt(2x) I'm supposed to solve it using u-substitution but im stumped

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{1}{9+\sqrt{2x}}~dx}\)

OpenStudy (anonymous):

yes

OpenStudy (freckles):

I dare you to try subbing the whole bottom as u

OpenStudy (anonymous):

that is du = 1/sqrt(2x) dx but what do i do about the sqrt(2x) ?

OpenStudy (freckles):

\[u=9+\sqrt{2x} \\ u-9=\sqrt{2x}\]

OpenStudy (freckles):

\[\int\limits \frac{u-9}{u} du\]

OpenStudy (anonymous):

oh ok that makes alot of sense

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

x - 9(ln|9 + sqrt(2x)|) + C

OpenStudy (freckles):

just in case... \[u=9+\sqrt{2x} \\ du=\frac{1}{\sqrt{2x}} dx \\ \text{ replace } \sqrt{2x} \text{with } u-9 \\ du=\frac{1}{u-9} dx \\ (u-9) du =dx\] ok now checking your final answer...

OpenStudy (freckles):

why did you replace the first u with x and not 9+sqrt(2x)?

OpenStudy (freckles):

should have u-9ln|u|+C where u=9+sqrt(2x)

OpenStudy (freckles):

that first u you put x?

OpenStudy (anonymous):

can't you split it up into u/u - 9/u ?

OpenStudy (freckles):

you can

OpenStudy (freckles):

but you are integrating with respect to u not one x and the other u

OpenStudy (anonymous):

oh ok so the 1 would integrate back into u which i would then replace

OpenStudy (freckles):

right because we are integrating with respect to u because of the du part

OpenStudy (freckles):

\[\int\limits (1-\frac{9}{u} ) du\\ u-9 \ln|u|+C\]

OpenStudy (freckles):

and final step is to replace any u you see with 9+sqrt(2x)

OpenStudy (anonymous):

ok thank you, I'll practice more problems like this one

OpenStudy (freckles):

k

OpenStudy (freckles):

I think there is another way without u sub to do this one would you like to see?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please

OpenStudy (freckles):

\[\frac{1}{9+\sqrt{2x}} \cdot 1 \\=\frac{1}{9+\sqrt{2x}} \cdot \frac{9-\sqrt{2x}}{9-\sqrt{2x}} \\ =\frac{9-\sqrt{2x}}{81-2x}\] oops spoke to soon.

OpenStudy (anonymous):

i actually tried doing it that way but i quickly figured out i would involve multiple substitutions which idk how to do yet

OpenStudy (freckles):

\[\frac{1}{9+\sqrt{2x}} =\frac{1}{\sqrt{2x}} \cdot \frac{\sqrt{2x}}{9+\sqrt{2x}} \\ =\frac{1}{\sqrt{2x}} \cdot \frac{\sqrt{2x}+9-9}{9+\sqrt{2x}} \\ =\frac{1}{\sqrt{2x}} \cdot (\frac{\sqrt{2x}+9}{9+\sqrt{2x}}-\frac{9}{9+\sqrt{2x}}) \\ =\frac{1}{\sqrt{2x}} (1-\frac{9}{9+\sqrt{2x}}) \\ =\frac{1}{\sqrt{2x}}- \frac{1}{\sqrt{2x}} \cdot \frac{9}{9+\sqrt{2x}}\] yea you still might prefer a substitution on the second one

OpenStudy (freckles):

so yep I spoke way too soon earlier my bad

OpenStudy (anonymous):

hahaha no problem but seriously thank you for the help

OpenStudy (freckles):

np